OpenStudy (anonymous):

What is the solution of this linear system? 4x+y=-1, y=-2x+3

OpenStudy (anonymous):

-4x - 1 = -2x + 3 -1 - 3 = -2x + 4x

OpenStudy (anonymous):

-4 = 2x -4/2 = 2x -2 = x.

OpenStudy (anonymous):

solve for y:

OpenStudy (anonymous):

y = -2x + 3 y = -2(-2) + 3 y = 4 + 3 y = 7.

OpenStudy (anonymous):

(-2, 7) are your coordinates. :)

OpenStudy (raffle_snaffle):

Matrix?

OpenStudy (raffle_snaffle):

You can solve by using a matrix